# GEARS - A Fully Run-Time Configurable Geant4 Application

### v1

@@ Abstract
The Geant4 toolkit is the standard for simulating the passage of particles through matter, but its conventional architecture often requires users to modify and recompile C++ code to alter fundamental simulation parameters such as geometry, physics list, and primary particle source. This architectural constraint introduces significant friction for new users and slows down the experimental iteration cycle. This paper introduces GEARS (Geant4 Example Application with Rich features yet Small footprint), a universally applicable Geant4 application that fundamentally addresses this issue. GEARS achieves complete simulation configurability without C++ recompilation by strictly utilizing external configuration methods: Geometry is defined via simple text-based configuration, the Physics List is selected via the standard PHYSLIST environment variable, and the Primary Source is defined through the General Particle Source (GPS) macro commands. Furthermore, regarding GEARS as an application instead of a framework, key features include a flat Ntuple structure with short variable names for highly efficient analysis via TTree::Draw() and a solution for capturing vital Step Zero data. Output creation is also fully managed via run-time macro commands and volume properties. The project is distributed as a ready-to-use Docker container to eliminate compilation barriers. Through these design considerations, GEARS transforms Geant4 into a practical, ready-to-use tool, enabling users to rapidly prototype and execute simulations for diverse experiments solely through simple text configuration files, without ever needing to modify or compile the underlying C++ source code.
@@ 10 Conclusion
  • Primary Particle Configurability: Exclusive reliance on the General Particle Source (GPS) and traditional /gun/ macros ensures the primary particle characteristics are dynamically configured via text files.

### v2

@@ Abstract
The Geant4 toolkit is the standard for simulating the passage of particles through matter, but its conventional architecture often requires users to modify and recompile C++ code to alter fundamental simulation parameters such as geometry, physics list, and primary particle source. This architectural constraint introduces significant friction for new users and slows down the experimental iteration cycle. This paper introduces GEARS (Geant4 Example Application with Rich features yet Small footprint), a universally applicable Geant4 application that fundamentally addresses this issue. GEARS achieves complete simulation configurability without C++ recompilation by strictly utilizing external configuration methods: Geometry is defined via simple text-based configuration, the Physics List is selected via the standard PHYSLIST environment variable, and the Primary Source is defined through the General Particle Source (GPS) macro commands. Furthermore, regarding GEARS as an application instead of a framework, key features include a flat Ntuple structure with short variable names for highly efficient analysis via TTree::Draw() and a solution for capturing vital initial step data. Output creation is also fully managed via run-time macro commands and volume properties. The project is distributed as a ready-to-use Docker container to eliminate compilation barriers. Through these design considerations, GEARS transforms Geant4 into a practical, ready-to-use tool, enabling users to rapidly prototype and execute simulations for diverse experiments solely through simple text configuration files, without ever needing to modify or compile the underlying C++ source code.
@@ 4.2 Plain Text Geometry
4.3 Performance of Text Geometry I/O
Using a text-based geometry file presents no performance penalty during the simulation run itself when compared to geometry defined directly in C++. This is because, during the Geant4 initialization phase, the text geometry parser translates the content of the external configuration file into the same underlying C++ objects used by manually coded definitions. While there is a minimal delay associated with reading and parsing the external text file during the application’s startup, this effect is negligible and virtually never noticed in the daily use of GEARS. Therefore, the choice of text geometry offers significant configurability benefits without sacrificing simulation speed.
4.4 Handling Complex Geometries
While the plain text geometry syntax is simple, it can be seamlessly used to define highly complex detectors. When geometry becomes exceedingly complex, the challenge shifts from defining the syntax to managing the large volume of component placement and parameter data. To handle this, users can leverage any high-level programming language that excels at string manipulation, such as Python or shell scripts, to programmatically generate the text geometry definition files. A short loop within one of these scripts can generate a text file containing thousands of lines of detector placements. The simple, non-hierarchical syntax of the text geometry is particularly suited for this kind of programmatic generation, allowing users to create sophisticated detector models without ever needing to touch or compile C++ code. This enables external scripting to handle the complexity, reinforcing GEARS’s core mission.
@@ 7.6 Dynamic Data Logging Control
7.7 Analysis Snippets
The combination of the flat ntuple data structure and the concise variable names is designed to maximize interactive analysis efficiency. The power of this design is best demonstrated through concrete examples of data querying using the ROOT TTree::Draw() command, which executes analysis logic directly on the data file without requiring a compiled analysis script. The following examples illustrate common analysis tasks performed directly within a ROOT session, where t is the pointer to the GEARS output TTree.
t->Draw(”x:y”, ”trk==1”, ”l”, 1, 0)
draws tracks of the primary particle (trk==1) on the $x$-$y$ plane. The "l" option forces the output to be drawn as lines, and the final arguments, 1, 0, instruct ROOT to process only the first entry (event) in the tree. It demonstrates that the GEARS output structure can be used not only for statistical analysis, but also for individual event visualization.
t->Draw(”pro”, ”trk>1 && stp==0”)
shows which physics process (pro) are involved in generating (stp==0) secondary particles (trk>1).
t->Draw(”x:y:z”, ”vlm==1”)
displays the three-dimensional spatial distribution of all step points (”hits”) within the detector volume whose copy number (vlm) is 1.
t->Draw(”pdg”)
provides a quick overview of the types of particles that were tracked.
t->Draw(”pro”, ”pdg==22 && stp!=0”)
plots physics processes related to $\gamma$-rays (pdg==22), but not the generation of $\gamma$-rays (stp!=0).
t->Draw(”x:y”, ”trk>1”)
plots the $x$-$y$ spatial distributions of secondary particles (trk>1) throughout the simulated geometry.
t->Draw(”de/dl:p”)
plots the energy loss per unit length (de/dl), which represents the stopping power, versus the particle’s momentum (p). This is a fundamental plot for validating energy deposition models.
t->Draw(”et[1]”)
draws the total energy spectrum recorded by the sensitive volume (detector) with copy number 1, stored in the dedicated energy tally variable et[1].
@@ 10 Conclusion
  • Primary Particle Configurability: Exclusive reliance on the General Particle Source (GPS) and traditional /gun macros ensures the primary particle characteristics are dynamically configured via text files.
@@ 11 Acknowledgements
12 Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work the author used Gemini in order to generate the initial draft of this manuscript. After using this tool, the author reviewed and edited the content as needed and takes full responsibility for the content of the published article.